OpenStudy (anonymous):

surface area of a shpere with a radius of 19

OpenStudy (anonymous):

i don't under stand???

OpenStudy (anonymous):

The formula for the surface area of a sphere is: [4\pi\r² Substitute and work out. Good luck!

OpenStudy (anonymous):

\[4\pi\] not [4/pi\ sorry.

OpenStudy (cwrw238):

4 * pi * r^2 = 4* pi * 19^2